# Parallel Program Analysis on Path Ranges

### v1

@@ 7.2 RQ 1: Ranged Program Analysis Using Two Identical Analyses
| overall | proof | alarm | add. | proof | alarm | |
| Se | ||||||
| Value | ||||||
| Pred | ||||||
| Bmc | ||||||
@@
Effectiveness. We report on the obtained results for RQ1 in the first two blocks of Tab. 1. Each row contains the number of overall correctly solved tasks, the number of correctly computed proofs and alarms, and the number of additionally correctly solved tasks. Herein, “additionally” refers to the basic analysis, i.e. Ra-Se has solved tasks that are not solved by Se. Moreover, Tab. 1 contains the number of incorrectly computed proofs and alarms.
When analyzing the effectiveness of ranged program analysis for symbolic execution, we observe that Ra-Se can solve in total tasks more than the basic symbolic execution, an increase of %. Moreover, there are tasks that can only be solved using ranged program analysis. For all additionally solved tasks, Ra-Se computed a valid alarm. The number of reported proofs is nearly constant, as Se and Ra-Se both have to check all paths in the program leading to a property violation for infeasibility.
For value analysis, we can also observe that Ra-Value can solve tasks not solved by value analysis and solves in total tasks more than the basic analysis, an increase of %. Again, the number of additionally solved tasks contains nearly exclusively alarms, as the value analysis is also a path-based analysis that does not apply abstraction mechanisms.
Having a look at predicate abstraction and BMC, we again notice that the use of ranged program analysis leads to uniquely solved tasks, namely for Ra-Pred and for Ra-Bmc. In contrast to value analysis and symbolic execution, the combination using either two instances of BMC or predicate abstraction within the ranged program analysis does not increase but slightly decreases the overall number of correctly solved tasks. This decrease is most likely caused by the fact that both analyses consider more than one path of the program, either through abstraction or full program encoding.
@@
Efficiency. For analyzing the efficiency of the ranged program analysis, we present in Fig. 7 scatter plots comparing the wall time taken by the tools in a log-scale, in case both tools can solve the task and a range is generated. On the x-axis, we depict the time taken by the basic analyses, on the y-axis the wall time of the ranged program analyses. A point in the scatter plot means that the basic analysis solved the tasks in seconds wall time and the ranged program analysis in seconds. For points below the diagonal, the ranged program analysis can solve the tasks faster than the basic analysis, otherwise, it is slower. The dashed lines indicate, where tools are twice as fast / twice as slow as the other. At the first glance, we realize that the ranged program analyses have a somewhat similar runtime to the basic analyses.
@@
For a more detailed comparison of the wall time taken, we use Fig. 8. We depict the time taken by Se on the x-axis, grouped in intervals. The height of each bar represents the median value of the time increase for ranged program analysis, where the width of each bar corresponds to the number of tasks that are solved by Se in the interval. For Se, we make two observations: Firstly, the ranged program analysis takes approximately twice as long to compute a solution for tasks that can be solved fast (in less than ten seconds) by the basic analysis. Secondly, we observe that for complex tasks the ranged program analysis becomes faster w.r.t. the basic analysis, such that we notice a comparable wall time of basic analysis and ranged program analysis for the complex tasks taking more than s to solve. Computing the test case to split the program into ranges takes a few seconds in nearly all cases. For tasks that can be solved by symbolic execution in less than ten seconds, this additional computational effort is the main factor for the time increase. As the time taken to compute a range is approximately the same for all tasks, it influences the overall wall time for complex tasks only to a small extent. For the more complex tasks, the advantages of ranged program analysis, namely analyzing different parts of the program in parallel, become visible. Unfortunately, the additional overhead caused by using ranged program analysis exceeds the advantages of parallelization, as the ranged program analysis is nearly as fast as the basic analysis for complex tasks, but not faster. We make the same observations for the other three basic analyses. The bar plots for the comparison can be found in Section A.
@@
Summary. Using the same analysis within ranged program analysis increases the effectiveness of the path-based analyses symbolic execution and value analysis. Each instance of the ranged program analyses can solve tasks that are not solved by the basic analyses used. The efficiency is comparable for larger and more complex tasks.
@@ 7.3 RQ 2: Using Different Analyses Within Ranged Program Analysis
Evaluation Plan. To investigate whether other analysis combinations benefit from a composition of ranged analyses, we evaluated three different compositions: We combine predicate abstraction with symbolic execution, value analysis, and BMC. Thereby, we get the three configurations Ra-Se-Pred, Ra-Value-Pred, and Ra-Bmc-Pred for ranged program analysis. We compare the ranged program analyses to the two basic analyses used and are again interested in effectiveness and efficiency.
@@
Effectiveness. To evaluate the effectiveness of the combinations using two different analyses within ranged program analysis, we have to compare the first and the third block of Table 1. The column add. now contains the tasks that neither of the two basic analyses can solve but the ranged program analysis, whereas the other columns are as explained in RQ 1.
@@
For Ra-Se-Pred, the ranged program analysis can solve in total tasks more than the symbolic execution can solve standalone. Unfortunately, the Ra-Se-Pred does not solve more tasks than the predicate abstraction can solve standalone. For most of these tasks, the predicate abstraction successfully verifies the assigned range, but the symbolic execution fails to solve its part (range) of the task. We can observe that the ranged program analysis can solve tasks that neither the symbolic execution nor the predicate abstraction can solve.
@@
We make the same observations for the other two combinations of value analysis with predicate abstraction and for BMC with predicate abstraction: Ra-Value-Pred has a higher effectiveness compared to value analysis, as the total number of solved tasks increases by . In contrast, predicate abstraction computes more correct answers than Ra-Value-Pred. The reason for the lower effectiveness compared to predicate abstraction is again the fact, that the value analysis fails to solve its assigned range in most cases. The ranged program analysis can solve tasks that neither predicate abstraction nor value analysis can solve.
@@
As BMC can solve more tasks than predicate abstraction on the benchmark, the combination Ra-Bmc-Pred within ranged program analysis solves in total tasks more than predicate abstraction, but fewer than BMC. Nevertheless, Ra-Bmc-Pred computes a correct solution for tasks, where both analyses do not succeed.
@@
Efficiency. We depict the scatter plots comparing the wall time of the basic analyses and the ranged program analysis in Fig. 9. For each ranged program analysis we present two scatter plots, once for a comparison with the first and once for the second basic analysis used. We obtain similar results as in RQ 1: The ranged program analysis needs approximately twice as long for simple tasks that can be solved within less than ten seconds and has a comparable overall run time for complex tasks. The bar plots depicting the comparison can be found in the Section B.
@@ 7.4 RQ 3: The Effect of Work Stealing
Evaluation Plan. Next, we investigate the effect of work stealing. Therefore, we enable work stealing for each of the three configurations of ranged program analysis used for RQ 2, now denoted Ws-Se-Pred, Ws-Value-Pred, and Ws-Bmc-Pred. We compare the results of the configurations using work stealing to the basic analyses and to ranged program analyses without work stealing and again compare effectiveness and efficiency.
Effectiveness. To evaluate the effectiveness of ranged program analysis applying work stealing and two different analyses, we have to compare the fourth block of Tab. 1 against the first and third blocks. The column add. in the fourth block now contains the tasks that neither of the two basic analyses can solve but the ranged program analysis using work stealing can, whereas the other columns are as explained in RQ 1.
First and foremost, we observe that work stealing has a huge positive effect. In comparison to the basic analyses, the ranged program analyses combining symbolic execution with predicate abstraction and BMC with predicate abstraction outperform both employed basic analyses when work stealing is enabled. Ws-Se-Pred can solve tasks more than symbolic execution and tasks more than predicate abstraction, an increase of % respectively %. In addition, Ws-Bmc-Pred solves tasks more than BMC and more than predicate abstraction, increasing the effectiveness by % respectively %. When combining value and predicate abstraction, work stealing also increases the effectiveness, although the positive effect is not as large as for the two other combinations, as predicate abstraction solves more tasks than Ws-Value-Pred. For the majority of these tasks, the splitter fails to generate a range and thus only the value analysis is executed on the full program. In comparison to the value analysis, Ws-Value-Pred is able to compute more proofs and alarms than the default analysis. The positive effect of work stealing also becomes visible when comparing the total number of correct answers computed by ranged program analyses that do and do not employ it, as depicted in Fig. 10. Each Venn diagram shown depicts on the left (in green) the number of tasks solved exclusively when using ranged program analysis without work stealing, on the right (in blue) the number of tasks solved exclusively when using work stealing and the intersection is labeled with the number of tasks solved by both. Using work stealing enables the ranged program analyses to solve additional tasks compared to Ra-Se-Pred, compared to Ra-Value-Pred and compared to Ra-Bmc-Pred.
@@ 7.5 RQ 4: Witness Joining in Ranged Program Analysis
When comparing the effectiveness of Ra-Value-Pred to predicate abstraction and value analysis (cf. Tab. 1), we notice that there are tasks only solved by the ranged program analysis, whereas are additional proofs. From these additional proofs are validated. The one proof is not validated by the validators, as they reach the memory limitations during validation.121212Due to the size of the generated witness a manual inspection is infeasible.
@@ 7.6 Threats to Validity
We have conducted the experiments using the SV-Benchmarks, the largest available dataset for C program verification. Although it is widely used, especially in the competition SV-Comp, our findings may not completely carry over to other real-world C programs.
It is unlikely that the implementation of the ranged program analysis suffers from bugs, as we obtain comparable results for the ranged program analysis in [24], where the ranged program analysis is realized using instrumentation instead of the CPA-based realization that we use in this paper. The additional incorrect answers computed for the Ra-Se are most likely not caused by an error in the ranged program analysis, as they are also observed in [24]. Moreover, we randomly selected some of the additionally raised incorrect alarms and verified, that the basic analyses also report the alarms, if the search space is limited to the path that was incorrectly reported by the ranged program analysis.
We did not redo the experiments from [14], where we compare the different splitting strategies for the configutations of ranged program analysis. In our prior experiments, Lb3 was the splitter the ranged program analyses achieved the best performance with. It might be the case that certain combinations of ranged program analyses perform even better when a different splitting strategy is used. In that case, the findings from our evaluation and conclusions drawn still remain valid.
The data collected may deviate when conducting a reproduction study due to a different experimental setup or environment. To account for small, expected measurement errors, we restrict the presentation of our data to two significant digits.
@@ 8 Related Work
Ranged program analysis runs different analyses on separate partitions of the search space, but accompanies this with work stealing. After the analyses finished, it joins their result. In the following, we discuss related work for all those aspects in individual subsections.

### v2

@@ 7.3 RQ 1: Ranged Program Analysis Using Two Identical Analyses
| overall | proof | alarm | par.only | proof | alarm | |
| Se | - | |||||
| Value | - | |||||
| Pred | - | |||||
| Bmc | - | |||||
@@
Effectiveness. We report on the obtained results for RQ1 in the first two blocks of Tab. 1. Each row contains the number of overall correctly solved tasks, the number of correctly computed proofs and alarms, and the number of tasks solved only by the parallel combinations but not by the basic analyses used in the configurations. Moreover, Tab. 1 contains the number of incorrectly computed proofs and alarms. When analyzing the effectiveness of ranged program analysis for symbolic execution, we observe that Ra-Se correctly solves in total tasks and Se in total tasks, meaning that Ra-Se can solve in total tasks more than the basic symbolic execution, an increase of %. Moreover, there are tasks that can be solved only using ranged program analysis but not Se. In all cases, Ra-Se computed a valid alarm. The number of reported proofs is nearly constant, as Se and Ra-Se both have to check all paths in the program leading to a property violation for infeasibility.
For value analysis, we see that Value solves in total tasks and Ra-Value tasks, meaning that Ra-Value solves in total tasks more than the basic analysis, an increase of %. Again, we observe that Ra-Value can solve tasks that are not solved by value analysis, nearly exclusively containing correctly raised alarms, as the value analysis is also a path-based analysis that does not apply abstraction mechanisms.
Having a look at predicate abstraction and BMC, we again notice that there are tasks that are solved by the ranged program analyses only and not by the basic analyses, namely for Ra-Pred and for Ra-Bmc. In contrast to value analysis and symbolic execution, the combination using either two instances of BMC or predicate abstraction within the ranged program analysis does not increase but slightly decreases the overall number of correctly solved tasks: Ra-Pred can solve overall tasks and the basic analysis Pred tasks, a decrease of . Similarly, Ra-Bmc solves tasks and Bmc tasks, a decrease of . This decrease is most likely caused by the fact that both analyses consider more than one path of the program, either through abstraction or full program encoding.
@@
For analyzing the efficiency of the ranged program analysis, we present in Fig. 7 scatter plots comparing the wall time taken by the tools in a log-scale, in case both tools can solve the task and a range is generated. On the x-axis, we depict the time taken by the basic analyses, on the y-axis the wall time of the ranged program analyses. A point in the scatter plot means that the basic analysis solved the tasks in seconds wall time and the ranged program analysis in seconds. For points below the diagonal, the ranged program analysis can solve the tasks faster than the basic analysis, otherwise, it is slower. The dashed lines indicate, where tools are twice as fast / twice as slow as the other. At first glance, we realize that the ranged program analyses have a somewhat similar runtime to the basic analyses.
@@
For a more detailed comparison of the wall time taken, we use Fig. 8. We depict the time taken by basic analyses on the x-axis, grouped in intervals. The height of each bar represents the median value of the time increase for ranged program analysis, where the width of each bar corresponds to the number of tasks that are solved by the basic analyses in the interval. For all four basic analyses, we make two observations: Firstly, the ranged program analysis takes approximately twice as long to compute a solution for tasks that can be solved fast (in less than ten seconds) by the basic analysis. Secondly, we observe that for complex tasks the ranged program analysis becomes faster w.r.t. the basic analysis, although there are some outliers, such that we notice a comparable wall time of basic analysis and ranged program analysis for the complex tasks taking more than s to solve.
To analyze why the ranged program analyses need more overall time to compute a solution compared to the basic analyses, especially for tasks that can be solved fast, we have a more detailed look at the individual steps: We analyze the time needed to compute the ranges. Therefore, we measured the time taken by Lb3 to compute a test case separately for all tasks and depicted the results as boxplots in Fig. 9. Therein, we depict the CPU time taken by Lb3 in Fig. 9a and the wall time in Fig. 9b. In the median, Lb3 needs seconds CPU time and overall seconds to compute the bounds, visualized using the orange lines. The boxes contain the median values for % resp. % of the tasks. Hence, % of all bounds are computed within seconds CPU and seconds wall time. The whisker contains % of all tasks, hence there are only very few cases where Lb3 needs more than resp. seconds.
As computing the test cases takes in the median seconds CPU time and seconds wall time, this additional computational effort is a main factor for the time increase. As the time taken to compute a range is approximately the same for all tasks, it influences the overall wall time for complex tasks only to a small extent. For the more complex tasks, the advantages of ranged program analysis, namely analyzing different parts of the program in parallel, become visible. Unfortunately, the additional overhead caused by using ranged program analysis exceeds the advantages of parallelization, as the ranged program analysis is nearly as fast as the basic analysis for complex tasks, but not faster.
@@
Summary. Using the same analysis within ranged program analysis increases the effectiveness of the path-based analyses symbolic execution and value analysis. Each instance of the ranged program analyses can solve tasks that are not solved by the basic analyses used. The efficiency is comparable for larger and more complex tasks, and computing the ranges yields a constant overhead.
@@ 7.4 RQ 2: Using Different Analyses Within Ranged Program Analysis
Evaluation Plan. To investigate whether other analysis combinations benefit from a composition of ranged analyses, we evaluated three different compositions: We combine predicate abstraction with symbolic execution, value analysis, and BMC. Thereby, we get the three configurations Ra-Se-Pred, Ra-Value-Pred, and Ra-Bmc-Pred for ranged program analysis. In the three configurations, we combine predicate abstraction, a technique that can effectively compute proofs with techniques that are good at detecting property violations. We compare the ranged program analyses to the two basic analyses used and are again interested in effectiveness and efficiency.
@@
Effectiveness. To evaluate the effectiveness of the combinations using two different analyses within ranged program analysis, we have to compare the first and the third block of Table 1. The column par.only now contains the tasks that neither of the two basic analyses can solve but the ranged program analysis, whereas the other columns are as explained in RQ 1.
@@
For Ra-Se-Pred, the ranged program analysis can solve in total tasks, that are tasks more than symbolic execution can solve standalone (). Unfortunately, Ra-Se-Pred does not solve more tasks than the predicate abstraction can solve standalone (). For most of these tasks, the predicate abstraction successfully verifies the assigned range, but the symbolic execution fails to solve its part (range) of the task. We can also observe that the ranged program analysis can solve tasks that neither the symbolic execution nor the predicate abstraction can solve.
@@
We make the same observations for the other two combinations of value analysis with predicate abstraction and for BMC with predicate abstraction: Ra-Value-Pred has a higher effectiveness compared to value analysis, as the total number of solved tasks increases by 281 (from to ). In contrast, predicate abstraction computes correct answers and thus more than Ra-Value-Pred. The reason for the lower effectiveness compared to predicate abstraction is again the fact, that the value analysis fails to solve its assigned range in most cases. The ranged program analysis can solve tasks that neither predicate abstraction nor value analysis can solve.
@@
BMC can solve more tasks than predicate abstraction on the benchmark ( resp. ) and the combination Ra-Bmc-Pred within ranged program analysis solves tasks. Hence, Ra-Bmc-Pred computes in total more correct answers than predicate abstraction, but fewer than BMC. Nevertheless, Ra-Bmc-Pred computes a correct solution for tasks, where both, predicate abstraction and BMC, do not succeed.
@@
Efficiency. We depict the scatter plots comparing the wall time of the basic analyses and the ranged program analysis in Fig. 10. For each ranged program analysis we present two scatter plots, one for a comparison with the first and one for the second basic analysis used. We obtain similar results as in RQ 1: The ranged program analysis needs approximately twice as long for simple tasks that can be solved within less than ten seconds and has a comparable overall run time for complex tasks. In Sec. 7.4, we compare the efficiency of ranged program analysis using two different basic analyses. In Fig. 11, we present a more detailed comparison of the overall run time of the basic analyses with the ranged program analysis.
@@ 7.5 RQ 3: The Effect of Work Stealing
Evaluation Plan. Next, we investigate the effect of work stealing. The idea of work stealing is motivated by the fact that the combination of different analyses within ranged program analyses has overall no positive effect on the effectiveness. We observe that these instances fail to solve tasks that one of the employed basic analyses can solve, as the other one does not complete the verification on the assigned range. Therefore, we enable work stealing for each of the three configurations of ranged program analysis used for RQ 2, now denoted Ws-Se-Pred, Ws-Value-Pred, and Ws-Bmc-Pred. We compare the results of the configurations using work stealing to the basic analyses and to ranged program analyses without work stealing and again compare effectiveness and efficiency.
Effectiveness. To evaluate the effectiveness of ranged program analysis applying work stealing and two different analyses, we have to compare the fourth block of Tab. 1 against the first and third blocks. The column par.only in the fourth block now contains the tasks that neither of the two basic analyses can solve but the ranged program analysis using work stealing can, whereas the other columns are as explained in RQ 1.
Work Stealing vs. Basic Analyses. First and foremost, we observe that work stealing has a huge positive effect. In comparison to the basic analyses, the ranged program analyses combining symbolic execution with predicate abstraction and BMC with predicate abstraction outperform both employed basic analyses when work stealing is enabled. Ws-Se-Pred can solve in total tasks, that are tasks more than symbolic execution (solves ) and tasks more than predicate abstraction (solves ), an increase of % respectively %. In addition, Ws-Bmc-Pred solves in total tasks, that are tasks more than BMC () and more than predicate abstraction (), increasing the effectiveness by % respectively %.
When combining value and predicate abstraction, work stealing also increases the effectiveness, although the positive effect is not as large as for the two other combinations: Ws-Value-Pred solves tasks, Value tasks, and Pred tasks. Hence, Ws-Value-Pred increases the number of solved tasks compared to Value by and thus the effectiveness, but solves tasks less compared to predicate analysis. For the majority of these tasks, the splitter fails to generate a range and thus only the value analysis is executed on the full program. In comparison to the value analysis, Ws-Value-Pred is able to compute more proofs and alarms than the default analysis.
Work Stealing vs. No Work Stealing. The positive effect of work stealing also becomes visible when comparing the total number of correct answers computed by ranged program analyses that do and do not employ it, as depicted in Fig. 12. Each Venn diagram shown depicts on the left (in green) the number of tasks solved exclusively when using ranged program analysis without work stealing, on the right (in blue) the number of tasks solved exclusively when using work stealing and the intersection is labeled with the number of tasks solved by both. Using work stealing enables the ranged program analyses to solve additional tasks compared to Ra-Se-Pred, compared to Ra-Value-Pred and compared to Ra-Bmc-Pred.
@@ 7.6 RQ 4: Witness Joining in Ranged Program Analysis
When comparing the effectiveness of Ra-Value-Pred to predicate abstraction and value analysis (cf. Tab. 1), we notice that there are tasks only solved by the ranged program analysis, whereas are additional proofs. From these additional proofs are validated. One additional proof is not validated by the validators, as they reach the memory limitations during validation.131313Due to the size of the generated witness a manual inspection is infeasible.
@@ 7.7 Threats to Validity
Internal Validity. It is unlikely that the implementation of the ranged program analysis suffers from bugs, as we obtain comparable results for the ranged program analysis in [26], where the ranged program analysis is realized using instrumentation instead of the CPA-based realization that we use in this paper. The additional incorrect answers computed for the Ra-Se are most likely not caused by an error in the ranged program analysis, as they are also observed in [26]. Moreover, we randomly selected some of the additionally raised incorrect alarms and verified, that the basic analyses also report the alarms, if the search space is limited to the path that was incorrectly reported by the ranged program analysis.
The data collected may deviate when conducting a reproduction study using the same experimental setup. To account for small, expected measurement errors, we restrict the presentation of our data to two significant digits.
External Validity. We have conducted the experiments using the SV-Benchmarks, one of the largest available datasets for C program verification. Although it is widely used, especially in the competition SV-Comp, our findings may not completely carry over to other real-world C programs.
In our experiments, we mainly focused on tasks from the ReachSafety category of the SV-Benchmarks. To assess the generalizability to other tasks, we additionally evaluated the best-performing configuration Ra-Bmc-Pred on the software-systems category from SV-Comp, which comprises of additional tasks. The tasks in the software-systems category are designed to be more representative of real-world use cases and include realistic implementations from the AWS C commons library or the Linux kernel. We observe that the employed configuration of BMC scales poorly to the software-systems tasks, especially since it performs worse than predicate abstraction in detecting property violations141414The results are contained in our artifact [25].. As a result, the combination of BMC and predicate abstraction within Ra-Bmc-Pred is deemed to be less effective compared to predicate abstraction. We still observe that work stealing has a positive effect when employed in ranged program analysis, as it increases the number of correctly solved tasks by around % compared to Ra-Bmc-Pred.
We did not redo the experiments from [15], where we compared the different splitting strategies for the configurations of ranged program analysis. In our prior experiments, Lb3 was the splitter the ranged program analyses achieved the best performance with. It might be the case that certain combinations of ranged program analyses perform even better when a different splitting strategy is used. In that case, the findings from our evaluation and conclusions drawn still remain valid.
The concept of ranged program analysis presented in this work is conceptually applicable to arbitrary program analyses. The range reduction presented in Sec. 3.2 can only be applied to analyses that are implemented in the framework of configurable program analysis. To overcome this limitation, we have presented a more generic approach based on program instrumentation in [26], that allows to use of arbitrary off-the-shelf program analyses. The instrumentation-based construction of ranged analyses can also be used in combination with work stealing. Combining correctness witnesses, work stealing, and instrumentation is planned as future work.
Data Availability Statement.
All experimental data and our open-source implementation are publicly available in our supplementary artifact [25] archived at Zenodo. Therein, we also included a virtual machine for reproducing the experiments.
@@ 8 Related Work
Ranged program analysis runs different analyses on separate partitions of the search space, but accompanies this with work stealing. After the analyses are finished, it joins their result. In the following, we discuss related work for all those aspects in individual subsections.